What sample size would we need in a follow up study if we want a margin error of 5% with 98% confidence?
Example 3: More Medicine Remember the example about the experiment with a certain medicine? we need in a follow-up study if we want a mar (Example 2) What sample size gin oferror of 5% with 98% confidence?
Example 2: New Medicine An experiment finds that 27%of53 subjects report improvement after using a new medicine I. Create a 95% confidence interval for the actual cure rate 2. Make it wider-90% confidence interval. 3. What are the advantages and the disadvantages? we at
Answer #1

The ME = Zcritical * SQRT(p * q/n).

Squaring and solving, n = (Zc/ME)2 * p * q

Here ME = 0.05, p = 0.27, q = 0.73

The Zc at \alpha = 0.02 is 2.33

Therefore n = (2.33/0.05)2 * 0.27 * 0.73 = 428.01

Therefore n = 429 (Rounding to the nearest Integer)